OpenStudy (anonymous):

2x^2 + 3x - 4 - x^2 +x +9

OpenStudy (anonymous):

solving for x?

OpenStudy (anonymous):

x^2+4 x+5

OpenStudy (anonymous):

(x+4)(x-1) = x(x-1) + 4(x-1) = x2 - x + 4x - 4 = x2 + 3x - 4

OpenStudy (anonymous):

Example: what are the factors of 6x2 − 2x = 0 ? 6 and 2 have a common factor of 2: 2(3x2 − x) = 0 And x2 and x have a common factor of x: 2x(3x − 1) = 0 And we have done it! The factors are 2x and 3x − 1, We can now also find the roots (where it equals zero): 2x will be 0 when x = 0 3x − 1 will be zero when x = 1/3 And this is the graph (see how it is zero at x=0 and x=1/3):